Circular No. 7974

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V4742 SAGITTARII H. W. Duerbeck, C. Sterken and J. N. Fu, Vrije Universiteit, Brussels,

report spectroscopy of V4742 Sgr with the Danish 1.54-m telescope (+ DFOSC) at the European Southern Observatory, La Silla, on Sept. 18.06 UT. The spectrum (range 350–900 nm) shows an Fe II nova shortly after maximum, with a strong continuum and moderately strong emission lines of H, Fe II (multiplets 42, 48, 49, 74), and Na I; the FWHM of Hα to Hδ is 2500 km/s. The H, Ca II, and Fe II lines have weak, blueshifted absorption components; the Fe II absorption is split, with sharp components at −1250 and −2100 and broad ones at −4000 km/s.

COMET C/2002 S1 (SKIFF) Additional astrometry, the following preliminary parabolic orbital ele-

ments, and an ephemeris appear on MPEC 2002-S31. This comet may be of short period.

T = 2002 May 16.687 TT  ω = 45.^{o}988
Ω = 355.013 }2000.0
q = 2.56581 AU  i = 27.916

COMET C/2002 O6 (SWAN) Schmidt photographs of C/2002 O6 taken by M. Jäger near Vienna on

Sept. 13.125 UT show no nuclear condensation and a 4′ × 1′ coma of total mag ∼ 9.5 with an 8′ tail in p.a. 0^{o}. K. Kadota, Ageo, Japan, reports (via S. Nakano, Sumoto, Japan) that his images taken on Sept. 10.809 and 17.813 also show no condensation, the coma having faded from m_{1} = 11.2 to 12.1 during the past week. Visual m_{1} estimates: Aug. 22.07, 6.2 (B. H. Granslo, Fjellhamar, Norway, 12×50 binoculars); 29.07, 7.6 (Granslo); Sept. 1.84, 8.1 (R. J. Bouma, Groningen, The Netherlands, 15×80 binoculars); 13.14, 10.0: (Bouma; twilight; altitude 11^{o}; difficult object).

COMET C/2002 O4 (HÖNIG) Visual m_{1} estimates: Aug. 11.99 UT, 7.9 (V. Znojil, Vrchtepla, Slo-

vak Rep., 12×60 binoculars); 18.04, 8.1 (N. Biver, Ablis, France, 7×50 binoculars); 31.90, 8.2 (B. H. Granslo, Fjellhamar, Norway, 12×50 binoculars); Sept. 4.88, 8.4 (R. J. Bouma, Groningen, The Netherlands, 15×80 binoculars); 13.82, 8.8 (A. Diepvens, Balen, Belgium, 0.15-m refractor).

2002 September 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green